# Diffusion Dynamics Models with Generative State Estimation for Cloth Manipulation

### v1

@@ A-A Simulation Setup
TABLE IV: Dataset statistics.
A-B Data Collection
To collect state estimation data, we set up a comprehensive multi-view system that incorporates up to four calibrated stereo-depth sensors, strategically placed at randomized viewing angles within predefined ranges. Cloth is initialized by given an randomly pick-and-place action. This configuration enables the generation of paired datasets consisting of fused point clouds alongside their corresponding ground-truth mesh states across multiple viewpoints. The system leverages SAPIEN’s advanced stereo depth simulation capabilities[45], which significantly reduces the sim-to-real gap by faithfully reproducing point cloud characteristics observed in real-world scenarios. This high-fidelity simulation approach ensures robust and reliable state estimation performance when transferred to physical environments. In our point cloud fusion process, we augment camera extrinsic parameters to simulate real-world calibration errors. Specifically, we introduce rotational variations ranging from $-1.5^{\circ}$ to $1.5^{\circ}$ and translational variations from -0.5 to 0.5 cm. To better mimic real-world conditions, we also simulate depth sensor noise and occlusion effects by applying random point dropout with ratios between 0.1 and 0.2, and introducing noise to the fused point cloud.
To collect dynamics data, we employ diverse action sampling strategies to generate a comprehensive dataset of 500K examples. Our sampling approach encompasses two key methodologies designed to capture realistic cloth manipulation scenarios. The first method involves applying directionally-randomized displacements to selected mesh vertices, with particular emphasis on folding-oriented actions where the cloth is manipulated to create various folding patterns. We also simulate picking and relocation actions by applying upward and translational movements to randomly selected vertices. The second methodology focuses on pair-wise vertex manipulation, where vertex pairs are selected based on their spatial distances to simulate actions such as folding one point of the cloth onto another. Each incremental action is precisely controlled, with magnitudes ranging from 0.02 to 0.05 units. To evaluate the model’s performance across different time horizons and assess the impact of auto-regressive inference error accumulation, we generate action sequences varying in length from 15 to 35 steps. All resultant mesh deformations throughout these sequences are meticulously recorded to capture the complete dynamics of the cloth’s behavior.
Appendix B Implementation Details
B-A Model Details
@@
TABLE V: Point cloud encoder hyperparameters.
@@
We design a transformer-based architecture for state estimation, which consists of a point cloud encoder, positional embedding module, and a series of transformer blocks. The model takes both point cloud observations and mesh states as input. The point cloud is first processed through a patch-based encoder, while the mesh states are embedded using a patchified positional encoding scheme. These features are then processed through transformer blocks with cross-attention mechanisms to predict the mesh state.
@@
Activation function GEGLU
@@
TABLE VI: Model hyperparameters.
@@
The Fourier feature mapping applies frequency-based encoding separately to each spatial dimension $(A_{x},A_{y},A_{z})$ of the action vectors using both sine and cosine functions, resulting in an intermediate representation of dimension $2\times 3\times F$, where $F$ is the number of Fourier frequencies. This representation is then processed through an MLP to produce the final action embeddings.
@@
TABLE VII: Action encoder hyperparameters.
B-B Training Details
@@
TABLE VIII: Training hyperparameters.
B-C Planning Details
For planning, we employ a hybrid approach combining Model Predictive Control (MPC) and Cross Entropy Method (CEM). Our planner optimizes action sequences by iteratively sampling actions, evaluating their outcomes using the learned dynamics model, and updating the sampling distribution based on the costs. To enhance planning efficiency, we introduce two key strategies: (1) an informed action sampling mechanism and (2) a grasp point selection method. For action sampling, we initialize the sampling distribution using a prior direction informed by the target state. Specifically, we identify the K vertices with highest mean squared error (MSE) between the current and target states, and compute a weighted average direction based on their distances to the grasp point:
@@ B-C Planning Details
where $s_{t}^{i}$ and $s_{c}^{i}$ represent the position of vertex $i$ in target and current states respectively, and $\tau$ is a temperature parameter that controls the concentration of the probability distribution. A lower temperature leads to more deterministic selection focusing on maximum displacement vertices, while a higher temperature enables more exploratory behavior. The grasp point is then sampled from this distribution:
@@
This probabilistic selection mechanism provides several advantages over deterministic maximum displacement selection: (1) it allows for exploration of different grasp points, (2) it can adapt to different manipulation scenarios by adjusting the temperature parameter, and (3) it provides a smoother transition between different grasp point candidates. The planning algorithm is outlined in Algorithm 1. Hyperparameters for model-based planning is listed in Table X.
B-D Baseline Implementation
We introduce details of baseline implementation.
GNNs
We adopt the implementation from [44]. We construct a comprehensive graph representation for modeling cloth dynamics, incorporating object particles, end-effector interactions, and material properties. The graph structure consists of four main components: (1) state and action representations, (2) particle attributes and instance information, (3) relation matrices for particle interactions, and (4) material-specific physics parameters. The state representation captures both spatial positions and temporal dynamics through a history buffer of $n_{his}$ frames and future predictions of $n_{future}$ frames. Each state vector contains the 3D positions $(x,y,z)$ of both cloth particles and the end-effector. We maintain a fixed-size particle set through Farthest Point Sampling (FPS) with an adaptive radius range of $[0.05,0.1]$. We show detailed parameters for graph construction below.
@@ B-D Baseline Implementation
TABLE IX: GNN model hyperparameters.
@@
1:Initial state $s_{i}$, target state $s_{t}$, dynamics model $f_{\theta}$, number of iterations $N$
2:Number of samples $K$, sequence length $L$, action bounds $[a_{min},a_{max}]$
@@
4:$a_{best}\leftarrow\text{None}$, $c_{best}\leftarrow\infty$
5:for $i=1$ to $N$ do
6:     $A_{mppi}\leftarrow$ SampleGaussian$(K/2,L,\mu,\sigma,[a_{min},a_{max}])$
7:     $A_{uniform}\leftarrow$ SampleUniform$(K/2,L,[a_{min},a_{max}])$
8:     $A\leftarrow$ Concatenate$(A_{mppi},A_{uniform})$
9:     $S_{pred}\leftarrow f_{\theta}(S,A)$triangleright Predict trajectories
10:     $C\leftarrow$ ComputeCost$(S_{pred},A,T)$triangleright Evaluate costs
@@
15:     $\mu,\sigma\leftarrow$ UpdateDistribution$(A,C,\tau)$triangleright Update using weighted averaging
16:     $\sigma\leftarrow\sigma\cdot(1-i/N)$triangleright Anneal exploration
@@
Parameter Value
Number of iterations 5
Samples per iteration 16
Sequence length 5
Action dimension 3
Initial std deviation 0.1
Temperature 1.0
TABLE X: Planning hyperparameters.
Appendix C Additional Results
Dynamics Prediction
Quantitative results on perception input scenario is presented in Table XI.We present more qualitative results on forward dynamics prediction in Figure 14 and Figure 15.
Type Method $\downarrow$ MSE $\downarrow$ CD $\downarrow$ EMD
($10^{-3}$) ($10^{-2}$) ($10^{-2}$)
T-shirt GNN 6.36 $\pm$ 1.30 8.88 $\pm$ 1.12 8.29 $\pm$ 1.94
Transformer 4.18 $\pm$ 0.73 4.26 $\pm$ 0.51 7.93 $\pm$ 0.70
DDM 0.55 $\pm$ 0.27 1.49 $\pm$ 0.13 3.22 $\pm$ 0.47
Cloth GNN 2.17 $\pm$ 1.44 5.02 $\pm$ 0.90 7.31 $\pm$ 4.65
Transformer 1.30 $\pm$ 0.65 2.27 $\pm$ 0.46 7.06 $\pm$ 2.08
DDM 0.66 $\pm$ 0.45 2.12 $\pm$ 0.54 5.51 $\pm$ 1.03
TABLE XI: Quantitative results of last frame dynamics prediction. Long-horizon dynamics prediction results with DPM perception noisy input. Errors represent a 95% confidence interval.
Figure 13: Analysis of failure modes. Systematic categorization and visualization of common failure cases, highlighting key limitations: (a) grasp point selection errors, (b) gripper-induced cloth damage, (c) dynamics prediction inaccuracies, and (d) complex garment features (e.g., zippers, down-filled materials).
Planning
We present more qualitative results in Figure 17 in the simulation environment on planning.
Failure Cases Analysis
We also present failure cases results in Figure 13. We observed four kinds of common cases that may cause failure of an action.
(a) Lift cloth
(b) Fold cloth
(c) Rotate cloth
(d) Pull cloth

### v2

@@ C.1 Data Collection
Table 6: Simulation physical parameters.
To collect state estimation data, we set up a comprehensive multi-view system that incorporates up to four calibrated stereo-depth sensors, strategically placed at randomized viewing angles within predefined ranges. Cloth is initialized given a randomly "pick-and-place" action. This configuration enables the generation of paired datasets consisting of fused point clouds alongside their corresponding ground-truth mesh states across multiple viewpoints. The system leverages SAPIEN’s advanced stereo depth simulation capabilities[48], which significantly reduces the sim-to-real gap by faithfully reproducing point cloud characteristics observed in real-world scenarios. This high-fidelity simulation approach ensures robust and reliable state estimation performance when transferred to physical environments. In our point cloud fusion process, we augment camera extrinsic parameters to simulate real-world calibration errors. Specifically, we introduce rotational variations ranging from $1.5^{\circ}$ to $1.5^{\circ}$ and translational variations from -0.5 to 0.5 cm. To better mimic real-world conditions, we also simulate depth sensor noise and occlusion effects by applying random point dropout with ratios between 0.1 and 0.2, and introducing noise to the fused point cloud.
To collect dynamic data, we employ diverse action sampling strategies to generate a comprehensive dataset of 500K examples. Our sampling approach encompasses two key methodologies designed to capture realistic cloth manipulation scenarios. The first method involves applying directionally-randomized displacements to selected mesh vertices, with particular emphasis on folding-oriented actions where the cloth is manipulated to create various folding patterns. We also simulate picking and relocation actions by applying upward and translational movements to randomly selected vertices. The second methodology focuses on pair-wise vertex manipulation, where vertex pairs are selected based on their spatial distances to simulate actions such as folding one point of the cloth onto another. Each incremental action is precisely controlled, with magnitudes ranging from 0.02 to 0.05 units. To evaluate the model’s performance across different time horizons and assess the impact of auto-regressive inference error accumulation, we generate action sequences varying in length from 15 to 35 steps. All resultant mesh deformations throughout these sequences are meticulously recorded to capture the complete dynamics of the cloth’s behavior.
C.2 Model Details
@@
Table 7: Point cloud encoder hyperparameters.
@@
We design a transformer-based architecture for state estimation, which consists of a point cloud encoder, a positional embedding module, and a series of transformer blocks. The model takes both point cloud observations and mesh states as input. In dynamics model, the model takes an additional input channel containing a binary mask, indicating grasped mesh vertices. The point cloud is first processed through a patch-based encoder, while the mesh states are embedded using a patchified positional encoding scheme. These features are then processed through transformer blocks with cross-attention mechanisms to predict the mesh state.
@@
Activation function GELU
@@
Table 8: Model hyperparameters.
@@
Table 9: Action encoder hyperparameters.
The Fourier feature mapping applies frequency-based encoding separately to each spatial dimension $(A_{x},A_{y},A_{z})$ of the action vectors using both sine and cosine functions, resulting in an intermediate representation of dimension $2\times 3\times F$, where $F$ is the number of Fourier frequencies. Given the input action $a\in\mathbb{R}^{B\times N\times 3}$, where $N$ is the number of actions and $3$ represents the dimension of $(x,y,z)$ coordinates, we compute the embedding $\mathbf{e}\in\mathbb{R}^{B\times N\times D_{3}}$ as:
where $D_{3}$ is the action embedding dimension, $d\in\{0,\ldots,D_{3}/6-1\}$, and $f_{d}=100^{d/(D_{3}/6)}$ are the Fourier feature frequencies. The resulting embedding $e$ provides a rich, high-dimensional representation of the action space. This representation is then processed through an MLP to produce the final action embeddings, which is later injected as the condition into our model through cross-attention layers.
C.3 Training Details
@@
Table 10: Training hyperparameters.
C.4 Planning Details
For planning, we employ a hybrid approach combining Model Predictive Control (MPC) and Cross Entropy Method (CEM). Our planner optimizes action sequences by iteratively sampling actions, evaluating their outcomes using the learned dynamics model, and updating the sampling distribution based on the costs. To enhance planning efficiency, we introduce two key strategies: (1) an informed action sampling mechanism and (2) a grasp point selection method. For action sampling, we initialize the sampling distribution using a prior direction informed by the target state. Specifically, we identify the K vertices with the highest mean squared error (MSE) between the current and target states, and compute a weighted average direction based on their distances to the grasp point:
@@ C.4 Planning Details
where $s_{t}^{i}$ and $s_{c}^{i}$ represent the position of vertex $i$ in target and current states, respectively, and $\tau$ is a temperature parameter that controls the concentration of the probability distribution. A lower temperature leads to more deterministic selection focusing on maximum displacement vertices, while a higher temperature enables more exploratory behavior. The grasp point is then sampled from this distribution:
@@
This probabilistic selection mechanism provides several advantages over deterministic maximum displacement selection: (1) it allows for exploration of different grasp points, (2) it can adapt to different manipulation scenarios by adjusting the temperature parameter, and (3) it provides a smoother transition between different grasp point candidates. The planning algorithm is outlined in Algorithm 1. Hyperparameters for model-based planning are listed in Table 11.
Parameter Value
Number of iterations 5
Samples per iteration 16
Sequence length 5
Action dimension 3
Initial std deviation 0.1
Temperature 1.0
Table 11: Planning hyperparameters.
@@
Table 12: GNN model hyperparameters.
C.5 State Estimation Baseline Implementation
To create the fairest possible comparison, we provided the GT canonical mesh to both GarmentNets and MEDOR. This isolates the evaluation to their performance for mapping a known shape to a deformed configuration in the observation space. We retrained the TRTM baseline from scratch on our data, and evaluated GarmentNets and MEDOR using their official pretrained checkpoints. The model input domain gap is minimal, as both our work and these baselines use the CLOTH3D dataset [49] with the same crumpled-state generation procedures. The evaluation is particularly fair for MEDOR for its test-time adaptation mechanism.
C.6 Dynamics Baseline Implementation
We introduce details of the dynamics baseline implementation.
GNNs
We adopt the implementation from [6]. We construct a comprehensive graph representation for modeling cloth dynamics, incorporating object particles, end-effector interactions, and material properties. The graph structure consists of four main components: (1) state and action representations, (2) particle attributes and instance information, (3) relation matrices for particle interactions, and (4) material-specific physics parameters. The state representation captures both spatial positions and temporal dynamics through a history buffer of $n_{his}$ frames and future predictions of $n_{future}$ frames. Each state vector contains the 3D positions $(x,y,z)$ of both cloth particles and the end effector. We maintain a fixed-size particle set through Farthest Point Sampling with an adaptive radius range of $[0.05,0.1]$. We show detailed parameters for graph construction in Table 12.
C.7 Manipulation Pipeline Details
@@
1:Initial state $s_{i}$, target state $s_{t}$, dynamics model $f_{\theta}$, number of iterations NN
2:Number of samples $K$, sequence length $L$, action bounds $a_{min},a_{max}]$
@@
[…]
5:for $i=1$ to NN do
[…]
9:  $S_{pred}\leftarrow f_{\theta}(S,A)$ $\triangleright$ Predict trajectories
[…]
@@
[…]
16:  $\sigma\leftarrow\sigma\cdot(1-i/N)$ $\triangleright$ Anneal exploration